OpenStudy (alyssajobug):

Many people believe Angel Oak is the oldest living thing in the United States east of the Rockies. This sprawling live oak has stood on John's Island for 1400 years. As is typical of the oldest oaks, its massive limbs touch the ground before curving upward, creating a majestic canopy that covers more than 17,000 square feet. A 220-year old oak tree produced acorns at two different rates. Until it reached maturity, it produced an average of 900 acorns a year. Once it reached maturity, it produced an average of 2200 acorns a year. Angel Oak has produced about 289,000 acorns in its lifetime. At what age did it reach maturity?

OpenStudy (nurali):

900 M + ( 220 - M) 2200 = 289000 900M + 484000 - 2200M = 289000 1300M = 195000 M = 195000 /1300 = 150 Years

OpenStudy (anonymous):

Let the age at maturity be M 900 M + ( 220 - M) 2200 = 289000 900M + 484000 - 2200M = 289000 1300M = 195000 M = 195000 /1300 = 150 YEARS ANSWER